# States Violating Both Locality and Noncontextuality Inequalities in Quantum Theory

### v2

@@ 2 CHSH inequality and KCBS inequality
$\left|\sum_{i=0}^{n-2}\alpha_{i}\alpha_{i+1}-\alpha_{n-1}\alpha_{0}\right|\leq n-2$ (6)
@@
$\left|\sum_{i=0}^{n-2}\langle X_{i}X_{i+1}\rangle-\langle X_{n-1}X_{0}\rangle\right|\leq n-2.$ (7)
@@
$\eqalign{\left|\left\langle\varphi\middle|\sum_{j=0}^{n-2}B_{j}B_{j+1}-B_{2m}B_{0}\middle|\varphi\right\rangle\right|&=4\sum_{j=0}^{n}\braket{\varphi}{\psi_{j}}\braket{\psi_{j}}{\varphi}-n\cr&=4n\frac{\cos(\pi/n)}{1+\cos(\pi/n)}-n\\ &>n-2.}$ (10)
@@ 3 States violating both CHSH inequality and $n$-noncontextuality inequality
$\left|\left\langle\varphi\middle|I\otimes\left(\sum_{j=0}^{n-2}B_{j}B_{j+1}-B_{n-1}B_{0}\right)\middle|\varphi\right\rangle\right|>n-2$ (15)
@@ 4 States violating both $6$-locality inequality and KCBS inequality
$\begin{array}[]{rl}&\left|\langle\varphi\middle|I\otimes(B_{0}B_{1}+B_{1}B_{2}+B_{2}B_{3}+B_{3}B_{4}-B_{4}B_{0})\middle|\varphi\rangle\right|\\[8.0pt] &=\left(\frac{1}{2}\right)^{2}\left(\frac{10}{1+\cos(\pi/5)}-5\right)+\left(\frac{\sqrt{3}}{2}\right)^{2}\left(\frac{20\cos(\pi/5)}{1+\cos(\pi/5)}-5\right)\\[8.0pt] &=\frac{5}{2}(\sqrt{5}-1)\\[8.0pt] &>3\end{array}$ (19)

### v3

@@ 2 CHSH inequality and KCBS inequality
$\begin{array}[]{rl}\left|\sum_{i=0}^{n-2}\alpha_{i}\alpha_{i+1}-\alpha_{n-1}\alpha_{0}\right|\leq n-2&(for\ even\ n)\\ \sum_{i=0}^{n-2}\alpha_{i}\alpha_{i+1}-\alpha_{n-1}\alpha_{0}\leq n-2&(for\ odd\ n)\end{array}$ (6)
@@
$\begin{array}[]{rl}\left|\sum_{i=0}^{n-2}\langle X_{i}X_{i+1}\rangle-\langle X_{n-1}X_{0}\rangle\right|\leq n-2&(for\ even\ n)\\ \sum_{i=0}^{n-2}\langle X_{i}X_{i+1}\rangle-\langle X_{n-1}X_{0}\rangle\leq n-2&(for\ odd\ n).\end{array}$ (7)
@@
$\eqalign{\left\langle\varphi\middle|\sum_{j=0}^{n-2}B_{j}B_{j+1}-B_{2m}B_{0}\middle|\varphi\right\rangle&=4\sum_{j=0}^{n}\braket{\varphi}{\psi_{j}}\braket{\psi_{j}}{\varphi}-n\cr&=4n\frac{\cos(\pi/n)}{1+\cos(\pi/n)}-n\\ &>n-2.}$ (10)
@@ 3 States violating both CHSH inequality and $n$-noncontextuality inequality
$\left\langle\varphi\middle|I\otimes\left(\sum_{j=0}^{n-2}B_{j}B_{j+1}-B_{n-1}B_{0}\right)\middle|\varphi\right\rangle>n-2$ (15)
@@ 4 States violating both $6$-locality inequality and KCBS inequality
$\begin{array}[]{rl}&\langle\varphi|I\otimes(B_{0}B_{1}+B_{1}B_{2}+B_{2}B_{3}+B_{3}B_{4}-B_{4}B_{0})|\varphi\rangle\\[8.0pt] &=\left(\frac{1}{2}\right)^{2}\left(\frac{10}{1+\cos(\pi/5)}-5\right)+\left(\frac{\sqrt{3}}{2}\right)^{2}\left(\frac{20\cos(\pi/5)}{1+\cos(\pi/5)}-5\right)\\[8.0pt] &=\frac{5}{2}(\sqrt{5}-1)\\[8.0pt] &>3\end{array}$ (19)